# Similarity to contraction semigroups:

### v1

@@ 4.1. Holbrook-type condition for semigroups
It is well-known that the equality (2.2) can be relaxed by replacing $R$ and $R^{-1}$ by any pair of bounded operators $A$ and $B$ whenever the families of operators $(T_{\gamma})_{\gamma\in\Gamma}$ and $(AT_{\gamma}B)_{\gamma\in\Gamma}$ are semigroup representations; see, for instance, [52, Theorem 5 & Corollary 1], [71, Proposition 4.2] or [4, Proposition 2.4] (where the latter result was proved in [28, Proposition 5.5.6]).
@@
Holbrook proves this by explicitly constructing an equivalent norm on $H$ under which $T$ becomes a contraction. This norm appears to be induced by an inner product, since it was shown to satisfy the parallelogram law. See also the proof of [67, Theorem 9.1] for a similar approach.
We will need the next extension of Holbrook’s result to the setting of representations of $G$ on $H$. Its proof relies on Proposition 4.1 and employs an appropriate quotient space to produce an equivalent norm making the representation contractive. Thus the Hilbertian nature of the norm becomes apparent. This approach simplifies the technical details, makes the argument more transparent, and reveals hidden aspects of Holbrook’s original construction [^1] It was claimed in [43, p. 164] (without proof) that the proof of (4.3) can be reduced to the case $T^{n}=AS^{n}B$ for $n\in\mathbb{Z}_{+}$.. Note that implicitly the quotient space appeared in the arguments leading to [5, Theorem 2.1.3 & Theorem 4.1] for $G=\mathbb{{Z}}_{+}$.
@@ 4.2. Holbrook-type condition away from the origin
Recall that if $A$ is a bounded operator on a Hilbert space such that $A^{N}$ is similar to a contraction for some $N\in\mathbb{{N}}$, then $A$ is similar to a contraction; see, for instance, [39, p. 912]. One may wonder whether a similar property holds for $C_{0}$-semigroups. More generally, given a semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ one may ask whether the subset of $0,\infty)$ consisting of those $t\in[0,\infty)$ such that $T(t)$ is similar to a contraction can be different from $0,\infty)$ and $\emptyset$. Somewhat surprisingly, Theorem 1.1 directly implies that the answer is negative, and we formalize it in the next corollary.
@@
Many results concerning similarity to contractions for single operators can be naturally extended to strongly continuous semigroups, provided that the semigroup consists of operators that are bounded from below. This key observation, which we establish in the proposition below, appears to have been overlooked in the literature. In particular, our result generalizes and strengthens Liapunov’s theorem for $C_{0}$-groups [36, Subsect. 7.2] and extends the main result of [79], where the authors proved that a uniformly continuous, quasi-compact, and bounded semigroup belongs to $\mathcal{SC}(H)$.
@@
see e.g. [62, Lemma 1.2.2] or [7, p. 487] (and cf. [35, Section 3]). Moreover, (4.17) is equivalent to the left invertibility of $T(t)$ for all $t\geq 0$, meaning that there exists $(S(t))_{t\geq 0}\subset\mathcal{L}(H)$ satisfying $S(t)T(t)=I$ for all $t\geq 0$. Moreover, $(S(t))_{t\geq 0}$ can be chosen to be a $C_{0}$-semigroup. See e.g. [7, Theorem 7.3] or [81, Theorem 1] for a relevant discussion.
@@
Proposition 4.11 implies a version of the Liapunov’s theorem for bounded from below (i.e. left-invertible) $C_{0}$-semigroups. It generalizes the result in [36, Corollary 7.2.5], where only exponentially stable $C_{0}$-groups were addressed.
@@
If $\mathcal{T}=(T(t))_{t\geq 0}$ is a $C_{0}$-semigroup with generator $E$, with $T(t^{\prime})$ bounded from below for some $t^{\prime}>0,$ then there exists $B\in\mathcal{L}(H)$ such that the semigroup generated by $E+B$ is similar to an isometric semigroup [81, Theorem 3] (see also [7, p. 497-498], so that $(T(t))_{t\geq 0}$ is similar to quasi-contraction semigroup. Thus, Proposition 4.11(i) and Corollary 4.12 are, in fact, corollaries of our main Theorem 1.4 below, whose proof is independent of both. However, it is instructive to formulate them here and provide with direct proofs as an illustration of condition (i) in Theorem 1.1.
@@ 4.3. Holbrook-type condition near the origin
Our arguments for the implication (iv) $\implies$ (i) in Theorem 1.2 will be based on the next result describing a splitting phenomenon for an evolution semigroup in $\mathcal{SQC}(L^{2}_{\nu}(\mathcal{H}))$, see Remark 4.14. Note [46, Proposition 8.9], where a similar statement was proved for a less demanding situation of the left shift on $\ell^{2}(\mathbb{N})$ with operator weight.
@@ 5. Similarity through quasi-similarity: putting the two sides together
In fact, we prove a more general statement by providing an upper bound on the similarity constant $C(\mathcal{T})$ of $\mathcal{T},$ which is of independent interest. While there are very few works addressing quantitative aspects of similarity constants, the importance of such bounds was emphasized and clarified, in particular, in [5], [44] and [71, Chapter 9].
@@ 6. Joint similarity constants in terms of individual ones
The next proposition links the individual similarity of $T(t),t\geq 0,$ to contractions to their joint similarity to contractions for all $t\geq 0$ via bounds on the individual similarity constants. It also provides a characterization of joint similarity to contractions in terms of the resolvent of the generator, which may be of interest for applications. The equivalence of statements (i) and (ii) is inspired by [44, Proposition 2.5], which concerns the lower semicontinuity of similarity constants with respect to the strong operator topology. It is worth noting that statement (ii) is local in nature, depending only on the similarity constants near zero.
@@ 7.1. Packel type semigroups and similarity to contractions
yields a $C_{0}$-semigroup. Following closely the proof of [65, property (I), pp. 242-243], we infer that $\mathcal{T}_{\mathbf{a}}$ is bounded on $L^{2}(\mathbb{{R}}_{+})\oplus L^{2}(\mathbb{{R}}_{+})$. Figure 2 illustrates the nature of $\mathcal{T}_{\mathbf{a}}.$
@@
In addition to Lemma 7.2, we will need a well-known asymptotic property of semigroup orbits stemming from [30] and already used in similar contexts. Given a $C_{0}$-semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ on a Hilbert space $H$, set
(7.10) $W(\mathcal{T}):=\{h\in H\,:\,\lim_{t\to\infty}T(t)h=0\mbox{ weakly}\}.$
Recall the next result proved in e.g. [65, Lemma 1], see also [29, Theorem 2.5] and [41, Theorems 4 and 5].
@@
Let $\mathcal{T}=(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Hilbert space $H$. If $\mathcal{T}\in\mathcal{SC}(H)$, then $W(\mathcal{T})\cap W(\mathcal{T}^{\ast})^{\perp}=\{0\}$.
The proof of Lemma 7.3 in [65] relies on the fact that the property $\mathcal{T}\in\mathcal{SC}(H)$ is equivalent to similarity of $\mathcal{T}$ to a semigroup admitting a unitary dilation, and the statement essentially reduces to the unitary case. The same argument applies if instead of the dilation one uses the factorization through a unitary group as in of Theorem 1.1(iii). Thus, in view of Theorem 1.1(iii) the following stronger version of the Lemma 7.3 holds.
@@
Let $\mathcal{T}=(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Hilbert space $H$. If there exists $\tau>0$ such that $T(\tau)$ is similar to a contraction, then $W(\mathcal{T})\cap W(\mathcal{T}^{\ast})^{\perp}=\{0\}$.
@@
Setting $W_{\mathbf{a}}(0)$ to be the zero operator and arguing as in the proof of [65, property (I), pp. 242-243], it is easy to see that
@@
will be called the Bhat-Skeide interpolating semigroup associated with $T$. A direct (but somewhat non-trivial) verification shows that $T(n)=I\otimes T^{n}$ for all $n\in\mathbb{Z}_{+},$ which explains our terminology. Moreover, $\mathcal{T}$ is contractive if and only if $T$ is a contraction, and $\mathcal{T}$ is bounded if and only if $T$ is power bounded. For justification of the above properties and further details, see [19, Section 2.1], and also the original result [10, Lemma 2.4] and the discussion in [74, Example 8.2]. Moreover, it is crucial that $\mathcal{T}$ belongs to $\mathcal{SC}\left(L^{2}(\mathbb{{T}})\otimes H\right)$ if and only if $T$ is similar to a contraction, see Proposition 7.5 below. Thus, the Bhat-Skeide interpolation provides a convenient method for constructing semigroup counterparts of operators that are not similar to contractions, avoiding the lengthy and technical computations found, for instance, in [65] and Section 7.1. This leads to the following observation, that is somewhat surprising.
@@
    Note that $(e^{E^{-1}t})_{t\geq 0}$ is a quasi-contraction $C_{0}$-semigroup (since its generator is bounded) that does not belong to $\mathcal{SC}(\mathcal{H})$. Indeed, if it belonged to $\mathcal{SC}(\mathcal{H})$, then $E^{-1}$ would be similar to a dissipative operator, which would imply that $E$ is similar to a dissipative operator, arriving at contradiction since $\mathcal{T}_{LM}\notin\mathcal{SC}(\mathcal{H})$; see also [36, Remark 5.5.4].
@@ 8.1. Similarity to contractions in terms of operator means and its interplay with control theory
Theorem 8.1 is a continuous counterpart of the discrete similarity criterion given in the essentially inaccessible preprint [24]. Its formulation and proof can be found in [70, Chapter 3]. The criterion states that $T\in\mathcal{L}(H)$ is similar to a contraction if and only if there are $\alpha,\beta>0$ and $C\in\mathcal{L}(H)$ such that
@@
Let $\overline{C}$ be the closure of $C,$ and let $\|\cdot\|_{\operatorname{dom}(\overline{C})}$ be the Hilbert space graph norm on $\operatorname{dom}(\overline{C})$, so that $\|h\|_{\operatorname{dom}(\overline{C})}^{2}=\|h\|^{2}+\|\overline{C}h\|_{K}^{2}$, $h\in\operatorname{dom}(\overline{C})$. Fix $h\in\operatorname{dom}(C)$ and let $\Omega$ be a null-set of $\mathbb{{R}}_{+}$ such that the set $\{CT(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ is separable in $K$. As the sets $\{T(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ and $\{CT(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ are separable in $H$ and $K$, respectively, it follows that $\{T(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ is a separable subset of $\left(\operatorname{dom}(\overline{C}),\|\cdot\|_{\operatorname{dom}(\overline{C})}\right)$. Hence, the closure of ${\operatorname{span}}\{T(t)h\,:t\in\mathbb{{R}}_{+}\setminus\Omega\}$ in $\left(\operatorname{dom}(\overline{C}),\|\cdot\|_{\operatorname{dom}(\overline{C})}\right)$ is a separable Hilbert space, and we denote it by $Y_{h}$. Consequently, the inclusion mapping $\mathcal{J}_{h}:(Y_{h},\|\cdot\|_{\operatorname{dom}(\overline{C})})\to(Y_{h},\|\cdot\|_{H})$ is a Borel isomorphism; see, for instance, [75, Corollary 4.5.5]. Since $\overline{C}:(\operatorname{dom}(\overline{C}),\|\cdot\|_{\operatorname{dom}(\overline{C})})\to K$ is continuous, it follows that $\overline{C}\mathcal{J}_{h}^{-1}:Y_{h}\to H$ is a Borel map, so the map $t\mapsto\overline{C}\mathcal{J}_{h}^{-1}T(t)h=CT(t)h$ from $\mathbb{{R}}_{+}\setminus\Omega$ to $K$ is Borel as well. Hence the latter map is Bochner measurable as $\{CT(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ is separable by hypothesis. Since $\Omega$ is a null-set of $\mathbb{{R}}_{+}$, we conclude that $F_{h}$ is Bochner measurable. ∎

### v2

@@ 4.1. Holbrook-type condition for semigroups
It is well-known that the equality (2.2) can be relaxed by replacing $R$ and $R^{-1}$ by any pair of bounded operators $A$ and $B$ whenever the families of operators $(T_{\gamma})_{\gamma\in\Gamma}$ and $(AT_{\gamma}B)_{\gamma\in\Gamma}$ are semigroup representations; see, for instance, [53, Theorem 5 & Corollary 1], [72, Proposition 4.2] or [4, Proposition 2.4] (where the latter result was proved in [29, Proposition 5.5.6]).
@@
Holbrook proves this by explicitly constructing an equivalent norm on $H$ under which $T$ becomes a contraction. This norm appears to be induced by an inner product, since it was shown to satisfy the parallelogram law. See also the proof of [68, Theorem 9.1] for a similar approach.
We will need the next extension of Holbrook’s result to the setting of representations of $G$ on $H$. Its proof relies on Proposition 4.1 and employs an appropriate quotient space to produce an equivalent norm making the representation contractive. Thus the Hilbertian nature of the norm becomes apparent. This approach simplifies the technical details, makes the argument more transparent, and reveals hidden aspects of Holbrook’s original construction [^1] It was claimed in [44, p. 164] (without proof) that the proof of (4.3) can be reduced to the case $T^{n}=AS^{n}B$ for $n\in\mathbb{Z}_{+}$.. Note that implicitly the quotient space appeared in the arguments leading to [5, Theorem 2.1.3 & Theorem 4.1] for $G=\mathbb{{Z}}_{+}$.
@@ 4.2. Holbrook-type condition away from the origin
Recall that if $A$ is a bounded operator on a Hilbert space such that $A^{N}$ is similar to a contraction for some $N\in\mathbb{{N}}$, then $A$ is similar to a contraction; see, for instance, [40, p. 912]. One may wonder whether a similar property holds for $C_{0}$-semigroups. More generally, given a semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ one may ask whether the subset of $0,\infty)$ consisting of those $t\in[0,\infty)$ such that $T(t)$ is similar to a contraction can be different from $0,\infty)$ and $\emptyset$. Somewhat surprisingly, Theorem 1.1 directly implies that the answer is negative, and we formalize it in the next corollary.
@@
Many results concerning similarity to contractions for single operators can be naturally extended to strongly continuous semigroups, provided that the semigroup consists of operators that are bounded from below. This key observation, which we establish in the proposition below, appears to have been overlooked in the literature. In particular, our result generalizes and strengthens Liapunov’s theorem for $C_{0}$-groups [37, Subsect. 7.2] and extends the main result of [80], where the authors proved that a uniformly continuous, quasi-compact, and bounded semigroup belongs to $\mathcal{SC}(H)$.
@@
see e.g. [63, Lemma 1.2.2] or [7, p. 487] (and cf. [36, Section 3]). Moreover, (4.17) is equivalent to the left invertibility of $T(t)$ for all $t\geq 0$, meaning that there exists $(S(t))_{t\geq 0}\subset\mathcal{L}(H)$ satisfying $S(t)T(t)=I$ for all $t\geq 0$. Moreover, $(S(t))_{t\geq 0}$ can be chosen to be a $C_{0}$-semigroup. See e.g. [7, Theorem 7.3] or [82, Theorem 1] for a relevant discussion.
@@
Proposition 4.11 implies a version of the Liapunov’s theorem for bounded from below (i.e. left-invertible) $C_{0}$-semigroups. It generalizes the result in [37, Corollary 7.2.5], where only exponentially stable $C_{0}$-groups were addressed.
@@
If $\mathcal{T}=(T(t))_{t\geq 0}$ is a $C_{0}$-semigroup with generator $E$, with $T(t^{\prime})$ bounded from below for some $t^{\prime}>0,$ then there exists $B\in\mathcal{L}(H)$ such that the semigroup generated by $E+B$ is similar to an isometric semigroup [82, Theorem 3] (see also [7, p. 497-498], so that $(T(t))_{t\geq 0}$ is similar to quasi-contraction semigroup. Thus, Proposition 4.11(i) and Corollary 4.12 are, in fact, corollaries of our main Theorem 1.4 below, whose proof is independent of both. However, it is instructive to formulate them here and provide with direct proofs as an illustration of condition (i) in Theorem 1.1.
@@ 4.3. Holbrook-type condition near the origin
Our arguments for the implication (iv) $\implies$ (i) in Theorem 1.2 will be based on the next result describing a splitting phenomenon for an evolution semigroup in $\mathcal{SQC}(L^{2}_{\nu}(\mathcal{H}))$, see Remark 4.14. Note [47, Proposition 8.9], where a similar statement was proved for a less demanding situation of the left shift on $\ell^{2}(\mathbb{N})$ with operator weight.
@@ 5. Similarity through quasi-similarity: putting the two sides together
In fact, we prove a more general statement by providing an upper bound on the similarity constant $C(\mathcal{T})$ of $\mathcal{T},$ which is of independent interest. While there are very few works addressing quantitative aspects of similarity constants, the importance of such bounds was emphasized and clarified, in particular, in [5], [45] and [72, Chapter 9].
@@ 6. Joint similarity constants in terms of individual ones
The next proposition links the individual similarity of $T(t),t\geq 0,$ to contractions to their joint similarity to contractions for all $t\geq 0$ via bounds on the individual similarity constants. It also provides a characterization of joint similarity to contractions in terms of the resolvent of the generator, which may be of interest for applications. The equivalence of statements (i) and (ii) is inspired by [45, Proposition 2.5], which concerns the lower semicontinuity of similarity constants with respect to the strong operator topology. It is worth noting that statement (ii) is local in nature, depending only on the similarity constants near zero.
@@ 7.1. Packel type semigroups and similarity to contractions
yields a $C_{0}$-semigroup. Following closely the proof of [66, property (I), pp. 242-243], we infer that $\mathcal{T}_{\mathbf{a}}$ is bounded on $L^{2}(\mathbb{{R}}_{+})\oplus L^{2}(\mathbb{{R}}_{+})$. Figure 2 illustrates the nature of $\mathcal{T}_{\mathbf{a}}.$
@@
In addition to Lemma 7.2, we will need a well-known asymptotic property of semigroup orbits, going back to [31] and already used in similar contexts. Given a $C_{0}$-semigroup $\mathcal{T}=(T(t))_{t\geq 0}$ on a Hilbert space $H$, set
$W(\mathcal{T}):=\{h\in H\,:\,\lim_{t\to\infty}T(t)h=0\;\text{weakly}\},$
and, given a bounded operator $T\in\mathcal{L}(H)$, define analogously
$W(T):=\{h\in H\,:\,\lim_{n\to\infty}T^{n}h=0\;\text{weakly}\}.$
Recall the following result proved in [66, Lemma 1], see also [30, Theorem 2.5] and [42, Theorems 4 and 5].
@@
Let $\mathcal{T}=(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Hilbert space $H$. If $\mathcal{T}\in\mathcal{SC}(H)$, then
$W(\mathcal{T})\cap W(\mathcal{T}^{\ast})^{\perp}=\{0\}.$
The proof of Lemma 7.3 in [66] relies on the fact that $\mathcal{T}\in\mathcal{SC}(H)$ is equivalent to similarity of $\mathcal{T}$ to a semigroup admitting a unitary dilation, and the statement then essentially reduces to the unitary case. We will need a different approach, relying on the discrete setting, to obtain the following stronger version. Its proof is comparatively direct and puts together several well-known facts from semigroup theory.
@@
Let $\mathcal{T}=(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Hilbert space $H$. If there exists $\tau>0$ such that $T(\tau)$ is similar to a contraction, then
$W(\mathcal{T})\cap W(\mathcal{T}^{\ast})^{\perp}=\{0\}.$
Proof.
Fix $\tau>0$ such that $T(\tau)$ is similar to a contraction, and let $R\in\mathcal{L}(H)$ be positive and invertible such that $S:=RT(\tau)R^{-1}$ is a contraction. It follows from [30, Theorem 3.1] that $W(S)=W(S^{\ast})$. Since $h\in W(T(\tau))$ if and only if
$\lim_{n\to\infty}S^{n}Rh=\lim_{n\to\infty}RT(\tau)^{n}h=0\quad\text{weakly},$
we infer that
(7.10) $W(T(\tau))=R^{-1}(W(S)).$
Similarly,
(7.11) $W(T(\tau)^{\ast})=R(W(S^{\ast}))=R(W(S)),$
since $S^{\ast}=R^{-1}T(\tau)^{\ast}R$. In addition, following the arguments given in [9, Lemma 3.3.1] (or [31, p. 788]), we have
$h\in R^{-1}(W(S)^{\perp})\;\iff\;Rh\in W(S)^{\perp}\;\iff\;h\in(R(W(S)))^{\perp},$
and using (7.11) we deduce that $R^{-1}(W(S)^{\perp})=(W(T(\tau)^{\ast}))^{\perp}$. Combining this with (7.10), we conclude that
(7.12) $W(T(\tau))\cap(W(T(\tau)^{\ast}))^{\perp}=\{0\}.$
Next we prove that
(7.13) $W(T(\tau))=W(\mathcal{T}).$
Clearly $W(\mathcal{T})\subseteq W(T(\tau))$. For the opposite inclusion, fix $h\in W(T(\tau))$. Since $\mathcal{T}$ is bounded, for any relatively compact subset $M\subset H$, we have
$\lim_{n\to\infty}\langle T(n\tau)h,f\rangle=0\quad\text{uniformly for }f\in M.$
For every $f\in H$, the set $M_{f}:=\{T(t)^{\ast}f\,:\,t\in[0,\tau)\}$ is relatively compact, so if $(t_{n})_{n\in\mathbb{{N}}}\subset[0,\infty)$ with $t_{n}\to\infty$, we may write $t_{n}=m_{n}\tau+r_{n}$ with $m_{n}\to\infty$, $m_{n}\in\mathbb{{N}}$ and $r_{n}\in[0,\tau)$. Then
$\langle T(t_{n})h,f\rangle=\langle T(m_{n}\tau)h,T(r_{n})^{\ast}f\rangle\to 0,$
since $T(r_{n})^{\ast}f\in M_{f}$. As $(t_{n})_{n\in\mathbb{N}}$ was arbitrary, this shows $\lim_{t\to\infty}T(t)h=0$ weakly, i.e. $h\in W(\mathcal{T})$. Thus (7.13) holds. An analogous argument yields $W(\mathcal{T}^{\ast})=W(T(\tau)^{\ast})$. Together with (7.12), this gives the claim. ∎
Remark 7.5.
Any invertible similarity $R$ suffices in the first part, its positivity is convenient but not essential. The property could be well-known, but we were not able to find a direct reference. For its global counterpart see e.g. [25, Theorem 3.4].
@@
Setting $W_{\mathbf{a}}(0)$ to be the zero operator and arguing as in the proof of [66, property (I), pp. 242-243], it is easy to see that
@@
will be called the Bhat-Skeide interpolating semigroup associated with $T$. A direct (but somewhat non-trivial) verification shows that $T(n)=I\otimes T^{n}$ for all $n\in\mathbb{Z}_{+},$ which explains our terminology. Moreover, $\mathcal{T}$ is contractive if and only if $T$ is a contraction, and $\mathcal{T}$ is bounded if and only if $T$ is power bounded. For justification of the above properties and further details, see [19, Section 2.1], and also the original result [10, Lemma 2.4] and the discussion in [75, Example 8.2]. Moreover, it is crucial that $\mathcal{T}$ belongs to $\mathcal{SC}\left(L^{2}(\mathbb{{T}})\otimes H\right)$ if and only if $T$ is similar to a contraction, see Proposition 7.6 below. Thus, the Bhat-Skeide interpolation provides a convenient method for constructing semigroup counterparts of operators that are not similar to contractions, avoiding the lengthy and technical computations found, for instance, in [66] and Section 7.1. This leads to the following observation, that is somewhat surprising.
@@
    Note that $(e^{E^{-1}t})_{t\geq 0}$ is a quasi-contraction $C_{0}$-semigroup (since its generator is bounded) that does not belong to $\mathcal{SC}(\mathcal{H})$. Indeed, if it belonged to $\mathcal{SC}(\mathcal{H})$, then $E^{-1}$ would be similar to a dissipative operator, which would imply that $E$ is similar to a dissipative operator, arriving at contradiction since $\mathcal{T}_{LM}\notin\mathcal{SC}(\mathcal{H})$; see also [37, Remark 5.5.4].
@@ 8.1. Similarity to contractions in terms of operator means and its interplay with control theory
Theorem 8.1 is a continuous counterpart of the discrete similarity criterion given in the essentially inaccessible preprint [24]. Its formulation and proof can be found in [71, Chapter 3]. The criterion states that $T\in\mathcal{L}(H)$ is similar to a contraction if and only if there are $\alpha,\beta>0$ and $C\in\mathcal{L}(H)$ such that
@@
Let $\overline{C}$ be the closure of $C,$ and let $\|\cdot\|_{\operatorname{dom}(\overline{C})}$ be the Hilbert space graph norm on $\operatorname{dom}(\overline{C})$, so that $\|h\|_{\operatorname{dom}(\overline{C})}^{2}=\|h\|^{2}+\|\overline{C}h\|_{K}^{2}$, $h\in\operatorname{dom}(\overline{C})$. Fix $h\in\operatorname{dom}(C)$ and let $\Omega$ be a null-set of $\mathbb{{R}}_{+}$ such that the set $\{CT(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ is separable in $K$. As the sets $\{T(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ and $\{CT(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ are separable in $H$ and $K$, respectively, it follows that $\{T(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ is a separable subset of $\left(\operatorname{dom}(\overline{C}),\|\cdot\|_{\operatorname{dom}(\overline{C})}\right)$. Hence, the closure of ${\operatorname{span}}\{T(t)h\,:t\in\mathbb{{R}}_{+}\setminus\Omega\}$ in $\left(\operatorname{dom}(\overline{C}),\|\cdot\|_{\operatorname{dom}(\overline{C})}\right)$ is a separable Hilbert space, and we denote it by $Y_{h}$. Consequently, the inclusion mapping $\mathcal{J}_{h}:(Y_{h},\|\cdot\|_{\operatorname{dom}(\overline{C})})\to(Y_{h},\|\cdot\|_{H})$ is a Borel isomorphism; see, for instance, [76, Corollary 4.5.5]. Since $\overline{C}:(\operatorname{dom}(\overline{C}),\|\cdot\|_{\operatorname{dom}(\overline{C})})\to K$ is continuous, it follows that $\overline{C}\mathcal{J}_{h}^{-1}:Y_{h}\to H$ is a Borel map, so the map $t\mapsto\overline{C}\mathcal{J}_{h}^{-1}T(t)h=CT(t)h$ from $\mathbb{{R}}_{+}\setminus\Omega$ to $K$ is Borel as well. Hence the latter map is Bochner measurable as $\{CT(t)h\,:\,t\in\mathbb{{R}}_{+}\setminus\Omega\}$ is separable by hypothesis. Since $\Omega$ is a null-set of $\mathbb{{R}}_{+}$, we conclude that $F_{h}$ is Bochner measurable. ∎